macroeconomics

Suppose that the labour force is currently 5 million and the unemployment rate is 8%. On August 1, government announces that it will create 40,000 new jobs to be filled over the next three months. Only people who were unemployed on August 1 will be hired. As a result, 130,000 people immediately apply for the jobs; half of these people had previously not been actively seeking work.

a. Has the labour force changed as a result of this announcement? 

multiple choice



b. The unemployment rate after the announcement but before the jobs are filled will be  %. Enter your responses rounded to one decimal place.
 
c. Once all the jobs are filled, the unemployment rate will be  %.
Answer #1

Correct answer:

a.

Yes, the labour force has changed as a result of this announcement

Reason: With half of the 110,000 people not actively seeking work (and thus out of labor force) apply for jobs, the labor force will increase by 110,000/2 = 55,000

b.

Unemployment rate = 0.12 = Unemployed/6,000,000

Unemployed = 720,000

After announcement labor force = 6,000,000 + 55,000 = 6,055,000

After announcement unemployment rate = 720,000/6,055,000 = 0.1189 or 11.89%

c.

After jobs are filled,

Unemployed = 720,000 - 48,000 = 672,000

Unemployment rate = 672,000/6,055,000 = 0.1109 or 11.09%


answered by: Zahidul Hossain